Let X be a ramdom variable with the following probability distribution


x: 1, 2, 3, 4


P(X=x): 0.4, 0.3, 0.2, 0.1


a. Find E(X) and Var(X)


b. is E(X) a possible value of X

1
Expert's answer
2022-02-20T09:19:30-0500

a. "E(X)=1*0.4+2*0.3+3*0.2+4*0.1=2."

"Var(X)=1^2*0.4+2^2*0.3+3^2*0.3+4^2*0.1-2^2=1.9."


b. E(X) is a possible value of X.